OpenStudy (anonymous):

how to calculate number of sheets sized 200mmx100mm to built a tank with 1400mm dia and 2000mm height

OpenStudy (ash2326):

Total area of sheet required = total surface area of the cylinder Total surface area= \( 2\pi r h+ 2\pi r^2 \) r= radius here diameter= 1400 mm so radius = 700 mm h= height= 2000 mm . so we have Total surface Area=\(\large{ 2\times 3.14\times 700\times 2000 + 2\times 3.14 \times 700^2 }\) we get surface area as= 11869200 mm^2 Area of one Sheet=200 \(\times\)100=20000 mm^2 So no. of sheets required= 11869200/20000=593.46 or 594 sheets